### v5

@@ 1.1 Our contributions
To utilize Theorem 1.3 to obtain our $O(\log\log n)$-approximate mechanism, our plan, roughly speaking, is to use Theorem 1.3 in the same fashion as Theorem 1.2: that is, we use a vector $d$ sampled from the distribution $\mathcal{D}$ in Theorem 1.3 as posted payments for the sellers. But implementing this plan takes some work in order to deal with the case where the cost-vector $c$ is not handled by Theorem 1.3 (i.e., we have $c(S^{*})>B/O(\log\log n)$). We cannot simply use the mechanisms of [NPS24, BCGL17], replacing Theorem 1.2 with Theorem 1.3, because these mechanisms use oracles, such as knapsack-cover oracle [NPS24], approximate XOS oracle [BCGL17], that are different from (and somewhat more-powerful than) a demand oracle. (In [NPS24], the set $S^{*}$ is obtained using an oracle they call a knapsack-cover oracle; whereas in [BCGL17], the case where $c(S^{*})$ is large is dealt with by assuming that one has an XOS oracle for an XOS-approximation of $v$.) We assume only demand-oracle access in Theorem 1.1, so we need to come up with some novel ideas to deal with the case where $c(S^{*})$ is large. Combining these with Theorem 1.3 leads to Theorem 1.1.
@@ 2 Preliminaries
A mechanism consists of an allocation rule $f$ along with a payment scheme $p$. In the setting of budget-feasible mechanism design, we are given as input the publicly-known information of the valuation function $v:2^{U}\to\mathbb{R}_{+}$, the budget $B\in\mathbb{R}_{+}$, and the seller’s reported costs $\{c_{e}\}_{e\in U}$. The allocation rule, or algorithm, $f$ outputs a subset $S\subseteq U$ of sellers whose items are purchased by the mechanism designer. We refer to the set $S$ returned by $f$ as the set of winners. The payment scheme $p$ then assigns to each seller a payment $p_{e}$. The utility of seller $e$, when her true private cost is $\overline{c}_{e}\geq 0$ is $u_{e}=p_{e}-\overline{c}_{e}$ if $e\in S$ and $u_{e}=p_{e}$ if $e\not\in S$. Each seller reports a cost that maximizes her own utility. Note that $f,\{p_{e}\}_{e\in U}$ are functions of $v,B$ and $c$. For ease of notation, we treat $v$ and $B$ as fixed and treat $f,\{p_{e}\}_{e\in U}$ as functions of $c$ only. Additionally, $u_{e}$ is a function of $v,B$, the reported costs $c$ and the true cost $\overline{c}_{e}$ of seller $e$. For ease of notation, we treat $v,B$ as fixed and refer to $u_{e}(\overline{c}_{e};c_{e},c_{-e})$ as the utility of seller $e$ when the reported costs are $c$ and the true cost of seller $e$ is $\overline{c}_{e}$. We seek a mechanism $\mathcal{M}=(f,p)$ satisfying the following properties.
@@
A budget feasible mechanism is a mechanism that satisfies the above properties. Our mechanism is randomized, and will satisfy the above properties with probability $1$, that is, under all realizations of its random bits. A randomized mechanism that is truthful with probability $1$ is also called a universally-truthful mechanism.
@@ 3 Finding a good distribution over threshold vectors
Let $\kappa\in(0,1)$ be such that $\operatorname{OPT_{LP}}(\kappa,S^{*})-\operatorname{OPT_{LP}}(\kappa^{2},S^{*})\geq\frac{v(S^{*})}{8\log\log n}$ as guaranteed by Lemma 3.3. The distribution $\mathcal{D}[S^{*}]$ is obtained by sampling a set $S\sim x$ where $x$ is the optimal solution for $\bigl(\operatorname{LP}{(\kappa,S^{*})}\bigr)$ and then taking the vector $d^{S,\kappa}$. By Lemma 3.2, we have that $\mathcal{D}[S^{*}]$ is a distribution of vectors in $\Delta^{S^{*}}(B)$. Additionally, note that $x$ can be computed in polynomial time, and also has polynomial sized support, since its the output of the Ellipsoid algorithm. Moreover, given $S$ and $\kappa$, the vector $d^{S,\kappa}$ can also be computed in polynomial using demand oracles. It follows that a complete description of $\mathcal{D}[S^{*}]$ (that is, a list of vectors in the support of $\mathcal{D}[S^{*}]$ and their probabilities) can be computed in polynomial time.
@@
Proof of Lemma 3.3.
Let $Z=\{2^{-2^{i}}:1\leq i\leq\log\log n\}$. We will find $\kappa\in Z$ satisfying $\operatorname{OPT_{LP}}(\kappa,S^{*})-\operatorname{OPT_{LP}}(\kappa^{2},S^{*})\geq\frac{1}{4\log\log n}v(S^{*})$. By a telescoping argument, we have,
@@ 4 The budget feasible mechanism
We now use the result in Section 3 to design a randomized budget feasible mechanism for subadditive valuations that achieves an approximation factor of $O(\log\log n)$, thereby proving Theorem 1.1. As mentioned in Section 1.2, we first randomly partition the set of sellers into two groups $U_{1},U_{2}$, and get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ and discard them. We then work over the sellers in $U_{2}$. We compute a demand set over $U_{2}$, with appropriately chosen prices. This results in a set $S^{*}\subseteq U_{2}$ that has good value compared to $\operatorname{OPT}$. When $c(S^{*})\leq\frac{B}{O(\log\log n)}$, we can utilize Theorem 3.1 to prove a good approximation ratio by sampling a vector $d\sim\mathcal{D}[S^{*}]$ and return $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. However, we also need to deal with the case when $c(S^{*})>\frac{B}{O(\log\log n)}$. In this scenario, we will show that one can construct a single vector $d^{\prime}$ that provides suitable thresholds. We show that a suitable choice for $d^{\prime}$ is to set $d^{\prime}_{e}$ to be a suitable scaling of $q_{e}$, where $q\in\mathbb{R}^{S^{*}}_{+}$ is an optimal solution to the LP: $\max\ q(S^{*})$ subject to $q(S)\leq v(S^{*})-v(S^{*}\setminus S)$ for all $S\subseteq S^{*}$. We select the set of elements $A$ for which $c_{e}\leq d^{\prime}_{e}$, and prune this set to construct a subset $R^{\prime}\subseteq A$ satisfying $\Omega(\operatorname{OPT})-v(e^{*})\leq v(R^{\prime})\leq O(\operatorname{OPT})$. We show that constructing $R^{\prime}$ this way admits budget-feasible threshold payments. Returning either $R^{\prime}$ or $e^{*}$ then gives a good approximation, since if $v(e^{*})$ is small then $v(R^{\prime})\geq\Omega(\operatorname{OPT})-v(e^{*})$ would be large. One caveat here is that we cannot really branch on whether $c(S^{*})\leq\frac{B}{O(\log\log n)}$ as this causes problems with truthfulness. So instead, we return one of $R$, $R^{\prime}$, or $e^{*}$, each chosen with a suitable probability. The algorithm is described in detail below. We show that Algorithm 1 is a distribution over monotone algorithms, and so the corresponding payments can be obtained using Theorem 2.3.
@@
Lemma 4.4.
Algorithm 1 can be combined with suitable payments to obtain a mechanism that is truthful, individually rational, has no positive transfers, and is budget feasible, with probability $1$.
@@
We argue that Algorithm 1 is monotone under all realizations of its random bits, which immediately implies by Theorem 2.3 that using threshold prices as payments yields truthfulness, individual rationality, and no positive transfers, with probability $1$.
@@
Now, suppose that the set returned is $R^{\prime}$. Let $e\in R^{\prime}$ and suppose that $e$ decreases their reported cost to $c^{\prime}_{e}<c_{e}$. As before, let $c=(c_{e},c_{-e})$ and $c^{\prime}=(c^{\prime}_{e},c_{-e})$. We claim that the sets $A$ and $R^{\prime}$ are the same under both inputs $c$ and $c^{\prime}$. As argued above, the set $S^{*}$ is the same under both inputs $c$ and $c^{\prime}$. This implies that the vector $q$ computed in step P is also the same under $c$ and $c^{\prime}$ (since the computation of $q$ is fixed once $S^{*}$ and $v$ are fixed). This then also implies that $A$ is the same under both inputs: since $e\in R^{\prime}\subseteq A$ under input $c$, we have $c^{\prime}_{e}<c_{e}\leq q_{e}\cdot\frac{4B}{V_{1}}$; for any other element $f\neq e$ belonging to $A$ under input $c$, we have $c^{\prime}_{f}=c_{f}\leq q_{f}\cdot\frac{4B}{V_{1}}$. Finally, since the computation of $R^{\prime}$ does not consider the costs, it follows that $R^{\prime}$ is also the same under both reported costs $c$ and $c^{\prime}$. In particular, since $e\in R^{\prime}$ under input $c$, we also have $e\in R^{\prime}$ under input $c^{\prime}$.
@@
Otherwise, the mechanism returns either $R$ or $R^{\prime}$. Note that we have argued above that if a winner $e$ changes her reported cost and remains a winner, then the set returned, as also the “intermediate” sets $S^{*}$, $A$, also do not change. This property, which is called no-bossiness in [NPS24], makes it quite convenient to reason about threshold payments.
@@
Note that the above discussion also shows that the threshold for a player $e$ can be computed in polytime. The threshold value $\tau^{\prime}_{e}$ for $e$ to belong to the set $S^{*}$ can be obtained using a demand oracle. If the set output is $R$, then the threshold for $e$ is $\min\{\tau^{\prime}_{e},d_{e}\}$. If the set output is $R^{\prime}$, then the threshold for $e$ to lie in $A$ is $\tau^{A}_{e}=\min\bigl\{\tau^{\prime}_{e},q_{e}\cdot\frac{4B}{V_{1}}\bigr\}$. Note that fixing $A$ also fixes the set $R^{\prime}$. Let $\overline{A}$ be the $A$-set obtained under input $(c^{\prime}_{e},c_{-e})$ when $e$ belongs to it, i.e., when $c^{\prime}_{e}<\tau^{A}_{e}$; let $\overline{R}$ be the corresponding $R^{\prime}$-set. Note $\overline{A}$ and $\overline{R}$ depend only on $c_{-e}$. One can infer that $\tau_{e}=\tau^{A}_{e}$, if $e\in\overline{R}$ and is 0 otherwise. ∎

### v6

@@ 1.1 Our contributions
To utilize Theorem 1.3 to obtain our $O(\log\log n)$-approximate mechanism, our plan, roughly speaking, is to use Theorem 1.3 in the same fashion as Theorem 1.2: that is, we use a vector $d$ sampled from the distribution $\mathcal{D}$ in Theorem 1.3 as posted payments for the sellers. But implementing this plan takes some work in order to deal with the case where the cost-vector $c$ is not handled by Theorem 1.3 (i.e., we have $c(S^{*})>B/O(\log\log n)$). We cannot simply use the mechanisms of [NPS24, BCGL17], replacing Theorem 1.2 with Theorem 1.3, because these mechanisms use oracles, such as knapsack-cover oracle [NPS24], approximate XOS oracle [BCGL17], that are different from (and somewhat more-powerful than) a demand oracle. (In [NPS24], the set $S^{*}$ is obtained using an oracle they call a knapsack-cover oracle; whereas in [BCGL17], the case where $c(S^{*})$ is large is dealt with by assuming that one has an XOS oracle for an XOS-approximation of $v$.) We assume only demand-oracle access in Theorem 1.1, so we need to come up with some novel ideas to deal with the case where $c(S^{*})$ is large. Combining these with Theorem 1.3 leads to Theorem $1.$
@@ 2 Preliminaries
A mechanism consists of an allocation rule $f$ along with a payment scheme pp. In the setting of budget-feasible mechanism design, we are given as input the publicly-known information of the valuation function $v:2^{U}\to\mathbb{R}_{+}$, the budget $B\in\mathbb{R}_{+}$, and the seller’s reported costs $\{c_{e}\}_{e\in U}$. The allocation rule, or algorithm, $f$ outputs a subset $S\subseteq U$ of sellers whose items are purchased by the mechanism designer. We refer to the set $S$ returned by $f$ as the set of winners. The payment scheme $p$ then assigns to each seller a payment $p_{e}$. The utility of seller $e$, when her true private cost is $\overline{c}_{e}\geq 0$ is $u_{e}=p_{e}-\overline{c}_{e}$ if $e\in S$ and $u_{e}=p_{e}$ if $e\not\in S$. Each seller reports a cost that maximizes her own utility. Note that $f,\{p_{e}\}_{e\in U}$ are functions of $v,B$ and $c$. For ease of notation, we treat $v$ and $B$ as fixed and treat $f,\{p_{e}\}_{e\in U}$ as functions of $c$ only. Additionally, $u_{e}$ is a function of $v,B$, the reported costs $c$ and the true cost $\overline{c}_{e}$ of seller $e$. For ease of notation, we treat $v,B$ as fixed and refer to $u_{e}(\overline{c}_{e};c_{e},c_{-e})$ as the utility of seller $e$ when the reported costs are $c$ and the true cost of seller $e$ is $\overline{c}_{e}$. We seek a mechanism $\mathcal{M}=(f,p)$ satisfying the following properties.
@@
A budget feasible mechanism is a mechanism that satisfies the above properties. Our mechanism is randomized, and will satisfy the above properties with probability 11, that is, under all realizations of its random bits. A randomized mechanism that is truthful with probability 11 is also called a universally-truthful mechanism.
@@ 3 Finding a good distribution over threshold vectors
Let $\kappa\in(0,1)$ be such that $\operatorname{OPT_{LP}}(\kappa,S^{*})-\operatorname{OPT_{LP}}(\kappa^{2},S^{*})\geq\frac{v(S^{*})}{8\log\log n}$ as guaranteed by Lemma $3.$ The distribution $\mathcal{D}[S^{*}]$ is obtained by sampling a set $S\sim x$ where $x$ is the optimal solution for $\bigl(\operatorname{LP}{(\kappa,S^{*})}\bigr)$ and then taking the vector $d^{S,\kappa}$. By Lemma 3.2, we have that $\mathcal{D}[S^{*}]$ is a distribution of vectors in $\Delta^{S^{*}}(B)$. Additionally, note that $x$ can be computed in polynomial time, and also has polynomial sized support, since its the output of the Ellipsoid algorithm. Moreover, given $S$ and $\kappa$, the vector $d^{S,\kappa}$ can also be computed in polynomial using demand oracles. It follows that a complete description of $\mathcal{D}[S^{*}]$ (that is, a list of vectors in the support of $\mathcal{D}[S^{*}]$ and their probabilities) can be computed in polynomial time.
@@
Proof of Lemma $3.$
Let $Z=\{2^{-2^{i}}:1\leq i\leq\log\log n\}$. We will find $\kappa\in Z$ satisfying $\operatorname{OPT_{LP}}(\kappa,S^{*})-\operatorname{OPT_{LP}}(\kappa^{2},S^{*})\geq\frac{1}{O(\log\log n)}v(S^{*})$. By a telescoping argument, we have,
@@ 4 The budget feasible mechanism
We now use the result in Section 3 to design a randomized budget feasible mechanism for subadditive valuations that achieves an approximation factor of $O(\log\log n)$, thereby proving Theorem $1.$ As mentioned in Section 1.2, we first randomly partition the set of sellers into two groups $U_{1},U_{2}$, and get an estimate of $\operatorname{OPT}$ from the sellers in $U_{1}$ and discard them. We then work over the sellers in $U_{2}$. We compute a demand set over $U_{2}$, with appropriately chosen prices. This results in a set $S^{*}\subseteq U_{2}$ that has good value compared to $\operatorname{OPT}$. When $c(S^{*})\leq\frac{B}{O(\log\log n)}$, we can utilize Theorem 3.1 to prove a good approximation ratio by sampling a vector $d\sim\mathcal{D}[S^{*}]$ and return $R=\{e\in S^{*}:c_{e}\leq d_{e}\}$. However, we also need to deal with the case when $c(S^{*})>\frac{B}{O(\log\log n)}$. In this scenario, we will show that one can construct a single vector $d^{\prime}$ that provides suitable thresholds. We show that a suitable choice for $d^{\prime}$ is to set $d^{\prime}_{e}$ to be a suitable scaling of $q_{e}$, where $q\in\mathbb{R}^{S^{*}}_{+}$ is an optimal solution to the LP: $\max\ q(S^{*})$ subject to $q(S)\leq v(S^{*})-v(S^{*}\setminus S)$ for all $S\subseteq S^{*}$. We select the set of elements AA for which $c_{e}\leq d^{\prime}_{e}$, and prune this set to construct a subset $R^{\prime}\subseteq A$ satisfying $\Omega(\operatorname{OPT})-v(e^{*})\leq v(R^{\prime})\leq O(\operatorname{OPT})$. We show that constructing $R^{\prime}$ this way admits budget-feasible threshold payments. Returning either $R^{\prime}$ or $e^{*}$ then gives a good approximation, since if $v(e^{*})$ is small then $v(R^{\prime})\geq\Omega(\operatorname{OPT})-v(e^{*})$ would be large. One caveat here is that we cannot really branch on whether $c(S^{*})\leq\frac{B}{O(\log\log n)}$ as this causes problems with truthfulness. So instead, we return one of $R$, $R^{\prime}$, or $e^{*}$, each chosen with a suitable probability. The algorithm is described in detail below. We show that Algorithm 1 is a distribution over monotone algorithms, and so the corresponding payments can be obtained using Theorem 2.3.
@@
Lemma $4.$
Algorithm 1 can be combined with suitable payments to obtain a mechanism that is truthful, individually rational, has no positive transfers, and is budget feasible, with probability 11.
@@
We argue that Algorithm 1 is monotone under all realizations of its random bits, which immediately implies by Theorem 2.3 that using threshold prices as payments yields truthfulness, individual rationality, and no positive transfers, with probability 11.
@@
Now, suppose that the set returned is $R^{\prime}$. Let $e\in R^{\prime}$ and suppose that $e$ decreases their reported cost to $c^{\prime}_{e}<c_{e}$. As before, let $c=(c_{e},c_{-e})$ and $c^{\prime}=(c^{\prime}_{e},c_{-e})$. We claim that the sets AA and $R^{\prime}$ are the same under both inputs $c$ and $c^{\prime}$. As argued above, the set $S^{*}$ is the same under both inputs $c$ and $c^{\prime}$. This implies that the vector $q$ computed in step P is also the same under $c$ and $c^{\prime}$ (since the computation of $q$ is fixed once $S^{*}$ and $v$ are fixed). This then also implies that AA is the same under both inputs: since $e\in R^{\prime}\subseteq A$ under input $c$, we have $c^{\prime}_{e}<c_{e}\leq q_{e}\cdot\frac{4B}{V_{1}}$; for any other element $f\neq e$ belonging to AA under input $c$, we have $c^{\prime}_{f}=c_{f}\leq q_{f}\cdot\frac{4B}{V_{1}}$. Finally, since the computation of $R^{\prime}$ does not consider the costs, it follows that $R^{\prime}$ is also the same under both reported costs $c$ and $c^{\prime}$. In particular, since $e\in R^{\prime}$ under input $c$, we also have $e\in R^{\prime}$ under input $c^{\prime}$.
@@
Otherwise, the mechanism returns either $R$ or $R^{\prime}$. Note that we have argued above that if a winner $e$ changes her reported cost and remains a winner, then the set returned, as also the “intermediate” sets $S^{*}$, AA, also do not change. This property, which is called no-bossiness in [NPS24], makes it quite convenient to reason about threshold payments.
@@
Note that the above discussion also shows that the threshold for a player $e$ can be computed in polytime. The threshold value $\tau^{\prime}_{e}$ for $e$ to belong to the set $S^{*}$ can be obtained using a demand oracle. If the set output is $R$, then the threshold for $e$ is $\min\{\tau^{\prime}_{e},d_{e}\}$. If the set output is $R^{\prime}$, then the threshold for $e$ to lie in AA is $\tau^{A}_{e}=\min\bigl\{\tau^{\prime}_{e},q_{e}\cdot\frac{4B}{V_{1}}\bigr\}$. Note that fixing AA also fixes the set $R^{\prime}$. Let $\overline{A}$ be the AA-set obtained under input $(c^{\prime}_{e},c_{-e})$ when $e$ belongs to it, i.e., when $c^{\prime}_{e}<\tau^{A}_{e}$; let $\overline{R}$ be the corresponding $R^{\prime}$-set. Note $\overline{A}$ and $\overline{R}$ depend only on $c_{-e}$. One can infer that $\tau_{e}=\tau^{A}_{e}$, if $e\in\overline{R}$ and is 0 otherwise. ∎